OpenStudy (anonymous):

How to multiply a square of a number starting with 5 without actual multiplication.

OpenStudy (anonymous):

If I'm reading this right you can use addition: \[5x^2=x+x+x+x+x+x+x+x+x+x\]

OpenStudy (anonymous):

Take 50 = 50+50+50+50...50 times

OpenStudy (anonymous):

oh i know maybe what you mean

OpenStudy (anonymous):

like \[45^2\]

OpenStudy (anonymous):

method is this: you will end in 25 one more than 4 is 5 5 times 4 is 2 answer is 2025

OpenStudy (anonymous):

or \[65^2\] one more than 6 is 7 6 times 7 is 42 answer 4225

OpenStudy (anonymous):

was that the question?

OpenStudy (anonymous):

@Satellite73...... if the no is 52 square....then.....what is the answer.....without actual multiplication....

OpenStudy (anonymous):

i dunno. i know the trick for b5^2 where b is a between 1 and 9. not for 52^2 but that can't be bad because you know (50+2)(50+2)=2500+200+4=2704

OpenStudy (anonymous):

of course all of these require some 'actual multiplication'

OpenStudy (anonymous):

52+52+52+..

OpenStudy (anonymous):

want a hoot, watch some of these http://www.youtube.com/watch?v=IIwlBjNLpjI&feature=related

OpenStudy (anonymous):

i love these. love them

OpenStudy (anonymous):

I am a genius!.... my brain is now faster than a calculator...

OpenStudy (anonymous):

Oh! I have to pay.....

OpenStudy (anonymous):

really? not on youtube

OpenStudy (anonymous):

isn't it? i love it!

OpenStudy (anonymous):

No for the Math fast tricks CD..

OpenStudy (anonymous):

oh, i should get it, but maybe not

OpenStudy (anonymous):

impressed with how long it took to get \[x=\frac{1\pm\sqrt{5}}{2}\] maybe

OpenStudy (anonymous):

Moderators all woke up, I see:-)